# Kummer-Artin-Schreier-Witt Theory

### v1

@@ 5.2 When $s>1$
In addition, Proposition 4.11 shows that (31) can be simplified to a minimal polynomial for $Y_{s-1}$ over $B_{s-1}$:
$\frac{1}{E_{s-1}(Y_{0},\ldots,Y_{s-3})+\lambda Y_{s-2}}\left(\frac{\lambda^{p-1}}{p}Y_{s-1}^{p}+Y_{s-1}+g_{s-1}(Y_{0},\ldots,Y_{s-1})\right)=X_{s-1}.$
@@ 6 The Group Structure
From Proposition 4.1 and Proposition 3.2, we have the (not necessarily algebraic) coordinate maps $\gamma:\mathcal{W}_{s}\rightarrow W_{s}$ and $\lambda:\mathcal{V}_{s}\rightarrow W_{s}$, given by
@@
respectively.
@@
Define a group structure on $\mathcal{W}_{s}$ (and similarly on $\mathcal{V}_{s}$) by the associated Witt vector group structure on $W_{s}$ via $\gamma$ (or via $\lambda$).
Recall also that $X_{i}^{\prime}\equiv X_{i}\pmod{\pi}$. This indicates the natural group structure on the special fibers, which coincides with the group structure on $W_{s}(\mathbb{F}_{p})$.
@@
We have a commutative diagram of group morphisms
@@
The commutativity of the first square follows from Definition 6.1.
The maps $\tilde{E}_{s,p}$ and $\tilde{\Psi}_{s}$ are defined by
$\begin{split}(Y_{0}^{\prime},\ldots,Y_{s}^{\prime})&\mapsto(E_{1,p}((Y_{0}^{\prime})),\ldots,E_{s+1,p}((Y_{0}^{\prime},\ldots,Y_{s}^{\prime}))),\text{ and}\\ (X_{0}^{\prime},\ldots,X_{s}^{\prime})&\mapsto(G_{1,p}((X_{0}^{\prime})),\ldots,G_{s+1,p}((X_{0}^{\prime},\ldots,X_{s}^{\prime}))),\end{split}$
respectively. Observe that each $E_{j,p}((Y_{0}^{\prime},\ldots,Y_{j-1}^{\prime}))$ (resp. $G_{j,p}((X_{0}^{\prime},\ldots,X_{j-1}^{\prime}))$) is a polynomial in $\mathbb{Z}_{(p)}[\zeta_{p^{j}}][Y_{0},\ldots,Y_{j-1}]$ (resp. in $\mathbb{Z}_{(p)}[\zeta_{p^{j}}][X_{0},\ldots,X_{j-1}]$). The commutativity of the second square follows from the identity
@@
Note that the composition $\mathcal{V}_{s}\rightarrow\mathbb{G}_{m}$ is precisely the morphism $\Psi_{s}((X_{0}^{\prime},\ldots,X_{s-1}^{\prime}))$. Additionally, $\tilde{E}_{s,p}\circ\gamma$ and $\tilde{\Psi}_{s}\circ\lambda$ are morphisms of algebraic group schemes, which are isomorphic on the generic fibers. This completes the proof of the commutative diagram’s structure. ∎
@@ 7 The Universality
In this section, we will establish the proof of Theorem 1.2. Our methodology follows standard procedures, as outlined in [11, Section 3] and [7, Section 6.4]. To begin, we will fix some notation. We define $\mathcal{W}_{1}:=\operatorname{Spec}R\left[X_{0},\frac{1}{1+\lambda X_{0}}\right]$. Let $R_{\lambda}:=R/\lambda$ and denote $\mathbb{G}_{m,R_{\lambda}}$ as the special fiber of $\mathbb{G}_{m}:=\operatorname{Spec}R\left[X,\frac{1}{X}\right]$. We have the natural closed immersion $\mathbb{G}_{m,R_{\lambda}}\hookrightarrow\mathbb{G}_{m}$ denoted by $\iota$. Regarding the computation of $\mathcal{W}_{1}$, the following exact sequence of sheaves on the étale site of $\operatorname{Spec}A$ is particularly useful (see [10, Lemma 1.1]):
@@
where $\alpha:x\mapsto 1+\lambda x$ and $\beta:t\mapsto t\mod\lambda$. The following result is derived from [11, Theorem 2.2].
@@
$0\to\mathcal{W}_{1}(B)\to\mathbb{G}_{m}(B)$ $=B^{*}\to\iota_{*}\mathbb{G}_{m,R_{\lambda}}(B)=(B/(\lambda B))^{*}\to$
@@
By Hilbert’s Theorem 90, we know that $\operatorname{H}^{1}(\operatorname{Spec}B,\mathbb{G}_{m})=0$. Given that the map $f:B^{*}\to(B/(\lambda B))^{*}$ is surjective, the exactness of the long exact sequence implies that $\operatorname{H}^{1}(\operatorname{Spec}B,\mathcal{W}_{1})\cong\operatorname{coker}(f)=0$. ∎

### v2

@@ 5.2 When $s>1$
In addition, Proposition 4.11 shows that (33) can be simplified to the minimal polynomial for $Y_{s-1}$ over $B_{s-1}$:
$\frac{1}{E_{s-1}(Y_{0},\ldots,Y_{s-3})+\lambda Y_{s-2}}\left(\frac{\lambda^{p-1}}{p}Y_{s-1}^{p}+Y_{s-1}+H_{s}(Y_{0},\ldots,Y_{s-1})\right)=X_{s-1}.$
@@ 6 The Group Structure
In this section, set $R:=\mathbb{Z}_{(p)}[\zeta_{p^{s}}]$, $\hat{R}:=\mathbb{Z}_{(p)}\llbracket\zeta_{p^{s}}-1\rrbracket$, and $\pi:=\zeta_{p^{s}}-1$. From Proposition 4.1 and Proposition 3.2, we have the (not necessarily algebraic) coordinate maps $\gamma:\mathcal{W}_{s}(R)\rightarrow W_{s}(\hat{R})$ and $\tau:\mathcal{V}_{s}(R)\rightarrow W_{s}(\hat{R})$, given by
@@
respectively. More precisely, each $X_{i}^{\prime}$ is computed inductively as follows:
$X_{0}^{\prime}$ $=\frac{1}{p\lambda}\ln(1+p\lambda X_{0}),$
$X_{1}^{\prime}$ $=\frac{1}{p\lambda}\ln\left(\frac{G_{2}(X_{0})+p\lambda X_{1}}{G_{2,p}((X_{0}^{\prime},0))}\right),$
$\vdots$
$X_{s-1}^{\prime}$ $=\frac{1}{p\lambda}\ln\left(\frac{G_{s}(X_{0},\ldots,X_{s-2})+p\lambda X_{s-1}}{G_{s,p}((X_{0}^{\prime},\ldots,X_{s-2}^{\prime},0))}\right).$
Conversely, given $(Y_{0}^{\prime},\ldots,Y_{s-1}^{\prime})\in W_{s}(\hat{R})$ (resp. $(X_{0}^{\prime},\ldots,X_{s-1}^{\prime})\in W_{s}(\hat{R})$), one can compute the corresponding point in $\mathcal{W}_{s}(\hat{R})$ (resp. $\mathcal{V}_{s}(\hat{R})$) as follows:
$Y_{0}$ $=\frac{1}{\lambda}\left(\exp\left(\lambda Y_{0}^{\prime}\right)-1\right),$
$Y_{1}$ $=\frac{1}{\lambda}\left(\exp\left(\lambda Y_{1}^{\prime}\right)E_{2,p}((Y_{0}^{\prime},0))-E_{2}(Y_{0})\right),$
$\vdots$
$Y_{s-1}$ $=\frac{1}{\lambda}\left(\exp\left(\lambda Y_{s-1}^{\prime}\right)E_{s,p}((Y_{0}^{\prime},\ldots,Y_{s-2}^{\prime},0))-E_{s}(Y_{0},\ldots,Y_{s-2})\right).$
(resp.
$X_{0}$ $=\frac{1}{p\lambda}\left(\exp\left(p\lambda X_{0}^{\prime}\right)-1\right),$ (36)
$X_{1}$ $=\frac{1}{p\lambda}\left(\exp\left(p\lambda X_{1}^{\prime}\right)G_{2,p}((X_{0}^{\prime},0))-G_{2}(X_{0})\right),$
$\vdots$
$X_{s-1}$ $=\frac{1}{p\lambda}\left(\exp\left(p\lambda X_{s-1}^{\prime}\right)G_{s,p}((X_{0}^{\prime},\ldots,X_{s-2}^{\prime},0))-G_{s}(X_{0},\ldots,X_{s-2})\right)$
). We can therefore naturally declare the group operations on $\mathcal{W}_{s}(\hat{R})$ and $\mathcal{V}_{s}(\hat{R})$.
@@
Define a group structure on $\mathcal{W}_{s}$ (and similarly on $\mathcal{V}_{s}$) over $\hat{R}$ by pulling back the Witt vector group structure on $W_{s}$ via the map $\gamma$ (or via $\tau$).
Remark 6.2.
Specifically, suppose $(X_{0,1},\ldots,X_{s-1,1})$ and $(X_{0,2},\ldots,X_{s-1,2})$ are points in $\mathcal{V}_{s}(\hat{R})$ with associated Witt vectors $(X_{0,1}^{\prime},\ldots,X_{s-1,1}^{\prime})$ and $(X_{0,2}^{\prime},\ldots,X_{s-1,2}^{\prime})$, respectively. Then we define their sum $(X_{0,3},\ldots,X_{s-1,3})$ as the unique point in $\mathcal{V}_{s}(\hat{R})$ satisfying
$\tau(X_{0,3},\ldots,X_{s-1,3})=(X_{0,1}^{\prime},\ldots,X_{s-1,1}^{\prime})+(X_{0,2}^{\prime},\ldots,X_{s-1,2}^{\prime}),$
where the right-hand side denotes the sum of Witt vectors in $W_{s}$.
@@
Definition 6.1 also makes $\psi_{s}:\mathcal{W}_{s}(R)\to\mathcal{V}_{s}(R)$ an isogeny.
Proof.
Suppose $(X_{0,1},\ldots,X_{s-1,1})$ and $(X_{0,2},\ldots,X_{s-1,2})$ are points in $\mathcal{V}_{s}$, as described in Remark 6.2, but now defined over $R$. Then the coordinates of the sum $(X_{0,3},\ldots,X_{s-1,3})$ are given by (36), and the sum $(X_{0,3}^{\prime},\ldots,X_{s-1,3}^{\prime})$ of the Witt vectors $(X_{0,1}^{\prime},\ldots,X_{s-1,1}^{\prime})$ and $(X_{0,2}^{\prime},\ldots,X_{s-1,2}^{\prime})$. Additionally, from the previous computation, we have
$\exp(p\lambda X_{j,3}^{\prime})=E_{j+1,p}(p(0,\ldots,0,X_{j,3}^{\prime}))=G_{j+1,p}((0,\ldots,0,X_{j,3}^{\prime})),$
so for $j\geq 1$, we have
$\exp\left(p\lambda X_{j,3}^{\prime}\right)G_{j+1,p}((X_{0,3}^{\prime},\ldots,X_{j-1,3}^{\prime},0))$ $=G_{j+1,p}((X_{0,3}^{\prime},\ldots,X_{j-1,3}^{\prime},X_{j,3}^{\prime}))$
$=G_{j+1,p}((X_{0,1}^{\prime},\ldots,X_{j,1}^{\prime}))\cdot G_{j+1,p}((X_{0,2}^{\prime},\ldots,X_{j,2}^{\prime}))$
$=\prod_{i=1}^{2}\left(G_{j+1}(X_{0,i},\ldots,X_{j-1,i})+p\lambda X_{j,i}\right).$
Applying the above identities to (36), we obtain:
$X_{0,3}$ $=\frac{1}{p\lambda}\left(\prod_{i=1}^{2}(1+p\lambda X_{0,i})-1\right),$
$X_{1,3}$ $=\frac{1}{p\lambda}\left(\prod_{i=1}^{2}(G_{2}(X_{0,i})+p\lambda X_{1,i})-G_{2}(X_{0,3})\right),$
$\vdots$
$X_{s-1,3}$ $=\frac{1}{p\lambda}\left(\prod_{i=1}^{2}\left(G_{s}(X_{0,i},\ldots,X_{s-2,i})+p\lambda X_{s-1,i}\right)-G_{s}(X_{0,3},\ldots,X_{s-2,3})\right).$
Hence, each $X_{j,3}\in R$ for $j=0,\ldots,s-1$. A similar argument applies to $\mathcal{W}_{s}$.
Recall also that $X_{i}^{\prime}\equiv X_{i}\pmod{\pi}$. This induces the natural group structure on the special fiber of $\mathcal{W}_{s}(R)\rightarrow\mathcal{V}_{s}(R)$, which coincides with the group structure on ~$\tilde{\wp}:W_{s}(\mathbb{F}_{p})\rightarrow W_{s}(\mathbb{F}_{p})$.
Finally, from the previous sections, if $\psi_{s}((Y_{0,1},\ldots,Y_{s-1,1}))=(X_{0,1},\ldots,X_{s-1,1})$, then $(Y_{0,1}^{\prime},\ldots\allowbreak\ldots,Y_{s-1,1}^{\prime})=(X_{0,1}^{\prime},\ldots,X_{s-1,1}^{\prime})$. The claim that $\psi_{s}:\mathcal{W}_{s}\rightarrow\mathcal{V}_{s}$ is a morphism of group schemes over $R$ follows immediately from the definition of group operations. ∎
Corollary 6.4.
With the notations of Corollary 4.13, the special fiber of $\psi_{s}:\mathcal{W}_{s}(R)\rightarrow\mathcal{V}_{s}(R)$ is the Artin-Schreier-Witt isogeny
$W_{s}(\mathbb{F}_{p})\rightarrow W_{s}(\mathbb{F}_{p}),\quad(y_{0},y_{1}+b_{1},\ldots,y_{s-1}+b_{s-1})\mapsto\wp(y_{0},y_{1}+b_{1},\ldots,y_{s-1}+b_{s-1}).$
Proposition 6.5.
We have a commutative diagram of group morphisms over $\hat{R}$:
@@
Furthermore, removing the map $W_{s}\rightarrow W_{s}$ yields a commutative diagram of homomorphisms of group schemes over $R$.
@@
The commutativity of the first square follows directly from Definition 6.1, which defines the group structure via the Witt vector group structure.
For the second square, we define the maps $\tilde{E}_{s,p}$ and $\tilde{\Psi}_{s}$ as follows:
$\begin{split}(Y_{0}^{\prime},\ldots,Y_{s}^{\prime})&\mapsto(E_{1,p}((Y_{0}^{\prime})),\ldots,E_{s+1,p}((Y_{0}^{\prime},\ldots,Y_{s}^{\prime}))),\\ (X_{0}^{\prime},\ldots,X_{s}^{\prime})&\mapsto(G_{1,p}((X_{0}^{\prime})),\ldots,G_{s+1,p}((X_{0}^{\prime},\ldots,X_{s}^{\prime}))).\end{split}$
The fact that these maps are group homomorphisms follows from (13) and (14). Additionally, each $E_{j,p}((Y_{0}^{\prime},\ldots,Y_{j-1}^{\prime}))$ (resp. $G_{j,p}((X_{0}^{\prime},\ldots,X_{j-1}^{\prime}))$) is a polynomial in $\mathbb{Z}_{(p)}[\zeta_{p^{j}}][Y_{0},\ldots,Y_{j-1}]$ (resp. $\mathbb{Z}_{(p)}[\zeta_{p^{j}}][X_{0},\ldots,X_{j-1}]$). Consequently, the compositions $\tilde{E}_{s,p}\circ\gamma$ and $\tilde{\Psi}_{s}\circ\tau$ define morphisms of algebraic group schemes over $R$ that are isomorphic on the generic fibers.
Recall that $\phi_{s}$ is defined in (9). The commutativity of the second square follows from:
@@
Finally, the commutativity of the third square has also been established in (9). This completes the proof of the commutativity of the entire diagram. ∎
@@ 7 The Universality
In this section, we establish the proof of Theorem 1.2. Our methodology follows standard procedures as outlined in [11, Section 3] and [7, Section 6.4]. To begin, we fix some notation. Let $R$ be a finite extension of $\mathbb{Z}_{(p)}[\zeta_{p^{s}}]$, and let $\pi$ be its uniformizer. Define $R_{\pi}:=R/\pi$, and denote by $\mathbb{G}_{m,R_{\pi}}$ the special fiber of $\mathbb{G}_{m}:=\operatorname{Spec}R\left[X,\frac{1}{X}\right]$. We have a natural closed immersion $\mathbb{G}_{m,R_{\pi}}\hookrightarrow\mathbb{G}_{m}$ denoted by $\iota$.
For the computation of $\mathcal{W}_{1}=\operatorname{Spec}R\left[Y_{0},\frac{1}{1+\lambda Y_{0}}\right]$, the following exact sequence of sheaves on the étale site of $\operatorname{Spec}R$ is particularly useful (see [10, Lemma 1.1]):
@@
where $\alpha:x\mapsto 1+\lambda x$ and $\beta:t\mapsto t\mod\pi$. The following result is derived from [11, Theorem 2.2].
@@
$0\to\mathcal{W}_{1}(B)\to\mathbb{G}_{m}(B)$ $=B^{*}\to\iota_{*}\mathbb{G}_{m,R_{\pi}}(B)=(B/(\pi B))^{*}\to$
@@
By Hilbert’s Theorem 90, we know that $\operatorname{H}^{1}(\operatorname{Spec}B,\mathbb{G}_{m})=0$. Given that the map $f:B^{*}\to(B/(\pi B))^{*}$ is surjective, the exactness of the long exact sequence implies that $\operatorname{H}^{1}(\operatorname{Spec}B,\mathcal{W}_{1})\cong\operatorname{coker}(f)=0$. ∎